OpenStudy (anonymous):

MEDAL AND FAN FOR WHO EVER CAN ANSWER MY QUESTION!!!! Janae distributed a bag of bird seed equally among 4 bird feeders in her backyard. When she finished, each bird feeder contained 6 ounces of bird seed. You can use the equation x/4 =6. Explain what x represents. How would you solve for x?

OpenStudy (anonymous):

X is the original amount of bird seed

OpenStudy (anonymous):

Which you distribute between bird feeders (this is represented by division) and this results in each getting 6

OpenStudy (anonymous):

To solve for x, just solve the equation x/4 = 6. An equation is like scales you can do the same stuff to both sides and it still balances. In this case multiply both sides by 4 to get x / 4 * 4 = 6 * 4 x = 24